OpenStudy (anonymous):

What does (1/2)^(2/3)^(3/4)^(4/5)...^(n/n+1) converge to?

OpenStudy (anonymous):

Is that an exponent for like 10 times? Or are those separate ?

OpenStudy (anonymous):

Infinite exponentation, somewhat like tetration

OpenStudy (mathmate):

Take log, then it is lim log((1/2)^(2/3)^(3/4)^(4/5)...^(n/n+1) ) =lim log(1/n+1) =zero e^0 =1 Therefore the limit is 1.

OpenStudy (anonymous):

no, that's not correct, if you evaluate the exponentation out to 15 or so it converges almost immediately on .6044...I want an exact answer.

OpenStudy (mathmate):

In multiplication, it is telescopic. Take log, then it is lim log((1/2)*(2/3)*(3/4)*(4/5)...*(n/n+1) ) =lim log(1/n+1) =zero e^0 =1 Therefore the limit is 1 when n->inf.

OpenStudy (anonymous):

you're not taking the limit though. think of it like an infinite sum or product, except exponentation.

OpenStudy (mathmate):

Does n to to infinity or is it finite?

OpenStudy (anonymous):

n goes to infinity. lim does not equal 1, try doing (1/2)^(2/3)^(3/4)^(4/5)^(5/6)^(6/7)^(7/8)^(8/9)^(9/10) and see that it almost immediately converges on .6044

OpenStudy (anonymous):

Usually convergences are integral or summation, but for this case it's neither...?

OpenStudy (anonymous):

why yes, you are correct.

OpenStudy (anonymous):

there are also infinite products, but this is an infinite exponentation.

OpenStudy (anonymous):

No idea. I'm used to seeing convergence problems that use integrals and summations, like taylor series or maclaurin series or P series... I can try thinking about it and see if I get anything. Is this calc 2?

OpenStudy (mathmate):

You're right, 1 is not the answer. It's a little more complicated. Will be back if I have new ideas.